Jefferson Lab HALL A Logbook for January 11 - All EXCEPT Automatic Entries
Main INDEX
Monday
346012 01/31/11 22:34 lselvy Change in Roc 8 boot script
346011 01/31/11 21:57 V. Sulkosky Re: helicity signals corrupted
346010 01/31/11 20:34 danez Left HRS Scaler Map
346009 01/31/11 19:25 lselvy FPP Status
346008 01/31/11 18:52 yez L-HRS: all HVs are on (not sure for VDCs)
346007 01/31/11 18:51 yez L-HRS: Start a long cosmic run overnight #1021
346006 01/31/11 18:24 yez L-HRS: can HVs of detectors be turned on now?
346005 01/31/11 18:18 lselvy HRS / HAND Coda Crash Resolved
346004 01/31/11 18:10 yez R-HRS cosmic ray long run started #20086
346003 01/31/11 18:04 yez R-HRS: Lemo input connector of Retiming module is bad
346002 01/31/11 16:22 nmuangma version 2 BigBite electronic
346001 01/31/11 15:29 yez LHRS-- powersupplies for crate #1 and #2 trip frequently
346000 01/31/11 15:10 orchen Runs with using 2,4,6,8,10,12 as trigger
345999 01/31/11 15:03 R. Michaels helicity signals corrupted
345998 01/31/11 14:44 Jack Right Arm bottom VDC high voltage
345997 01/31/11 13:44 V. Sulkosky BigBite Wire Chamber HV Off
345996 01/31/11 13:37 yez S1L ADC was not connected to Patch Panel but to Scope. It works fine
345995 01/31/11 13:21 orchen HAND runs for N4 gain calibration
345994 01/31/11 13:08 yez L-HRS retiming is done and start to take cosmic ray data
345993 01/31/11 12:22 kalyan lumi signals on RHRS
345992 01/31/11 12:02 R. Michaels R-HRS scaler repair, ready to adjust scaler.map
345991 01/31/11 11:07 solvignon R-arm S2 L-PMT ADC 0
345990 01/31/11 11:04 R. Michaels more on L-HRS DAQ (stack in, DAQ works)
345989 01/31/11 10:30 yez R-HRS Cosmic Ray run Update -- S1,S2m,Cer,VDC are all in good status
345988 01/31/11 10:08 David Anez Left HRS Electronics Update
345987 01/31/11 04:02 nmuangma BigBite Update
Sunday
345986 01/30/11 16:20 V. Sulkosky R-HRS VDC HV On
345985 01/30/11 15:58 nmuangma 1190 TDC readout with ThaScintPlane Class
345984 01/30/11 14:10 V. Sulkosky Re: N4-7L & N4-12L : Triping at low voltage
345983 01/30/11 11:38 orchen more info
345982 01/30/11 11:30 orchen HAND run for N4 gain
345981 01/30/11 11:26 orchen N4-7L & N4-12L : Triping at low voltage
Saturday
345980 01/29/11 21:46 orchen BigBite
345979 01/29/11 21:45 igorko Power stability problems
345978 01/29/11 19:23 Fansler Cryo Call In
345977 01/29/11 18:48 igorko cosmic run for HAND
345976 01/29/11 17:49 igorko cosmic run for HAND
345974 01/29/11 16:25 David Anez Right HRS Electronics Corrections
345973 01/29/11 16:21 nmuangma all noise/missing channels are fixed :version2
345972 01/29/11 16:17 nmuangma persistance missing channel in MWDC
345971 01/29/11 15:19 nmuangma The MWDC is REALLY sensitive, report if you are working anything related to it.
345970 01/29/11 15:05 nmuangma Using the Bottom Power Supply instead for MWDC Theshole. The top one needs to be fixed.
345969 01/29/11 14:41 nmuangma Problem with The Theshole power supply (front chamber), MWDC HV off
345968 01/29/11 14:22 nmuangma Tools must be put back in place!!! serious.
345967 01/29/11 08:44 orchen Weldment Power
Friday
345966 01/28/11 23:22 lselvy Changelog Setup
345965 01/28/11 20:02 V. Sulkosky Re: Scalers in trigger supervisor crate
345964 01/28/11 20:00 V. Sulkosky/ Bubbler for MWDC Moved
345963 01/28/11 17:59 nmuangma Done: t0 MWDC calibration, new 20110128 DB folder created.
345962 01/28/11 16:41 nmuangma TDC 1190 is not yet work.
345961 01/28/11 16:27 Jack Left Detector Stack is back in for the weekend
345960 01/28/11 16:26 nmuangma database change for mwdc.
345959 01/28/11 15:08 saw Scalers in trigger supervisor crate
345958 01/28/11 13:02 yez readjust retiming and start a new cosmic run #20065
345957 01/28/11 01:21 BigBite Preparation Update: all noise/missing channels are fixed.
Thursday
345956 01/27/11 13:32 yez readjust retiming and start a new cosmic run
345955 01/27/11 08:08 R. Michaels what to do when DAQ hangs
345954 01/27/11 03:01 nmuangma BigBite Preparation Update
345953 01/27/11 02:25 nmuangma HAND DAQ hangs
Wednesday
345952 01/26/11 23:44 nmuangma BigBite ADC spectrum (Good)
345951 01/26/11 19:31 yez RHRS single arm trigger is done and I will take some data overnight (do not kill the DAQ!)
345950 01/26/11 17:31 nmuangma BigBite ADC Spectrum
345949 01/26/11 15:16 R. Michaels CODA on Bigbite, Hand, and 2 HRS.
345948 01/26/11 15:12 yez RHRS single arm Trigger map
345947 01/26/11 15:07 yez RHRS single arm Trigger map
345946 01/26/11 15:07 yez RHRS single arm Trigger map
345945 01/26/11 15:06 yez RHRS single arm Trigger map
345944 01/26/11 15:06 yez RHRS single arm Trigger map
345943 01/26/11 15:06 yez RHRS single arm Trigger map
345942 01/26/11 15:06 yez RHRS single arm Trigger map
345941 01/26/11 14:53 nmuangma BigBite Electronic Mapping
345940 01/26/11 14:03 A. Camsonne Compton electron detector moved out from the Hall - Compton line leaked checked
345939 01/26/11 13:45 nmuangma BigBite Preparation Update
345938 01/26/11 11:59 Jack Left S0 removed from the stack
Tuesday
345937 01/25/11 19:54 yez RHRS HV crate replaced but does not function smoothly
345936 01/25/11 16:32 saw F1TDC configuration for bbvme1
345935 01/25/11 15:07 yez RHRS DAQ is very close to take comic ray data
345934 01/25/11 11:10 A. Camsonne RT cabling left arm
Monday
345933 01/24/11 15:30 R. Michaels HRS DAQ ready (sorta)
345932 01/24/11 14:24 atrig Trigger-Download
345931 01/24/11 13:13 fansler Lead Flows
Sunday
345930 01/23/11 14:45 Folts ESR crash call in
345929 01/23/11 11:43 Jack Port servers on Left Detector Stack
Saturday
345928 01/22/11 18:21 David Anez HRS Electronics Update
Friday
345927 01/21/11 15:56 Fansler RD1 N2
345926 01/21/11 12:55 R. Michaels CODA 2.6 instructions (I)
345925 01/21/11 10:51 fansler Magnet CRYO
345924 01/21/11 10:42 Fansler Hall A Cryo
Thursday
345923 01/20/11 16:01 atrig Trigger-Download
345922 01/20/11 16:00 atrig Trigger-Download
Wednesday
345921 01/19/11 15:29 atrig Trigger-Download
345920 01/19/11 15:28 atrig Trigger-Download
345919 01/19/11 15:26 atrig Trigger-Download
345918 01/19/11 15:25 atrig Trigger-Download
345917 01/19/11 15:21 atrig Trigger-Download
345916 01/19/11 15:16 atrig Trigger-Download
345915 01/19/11 12:04 atrig Trigger-Download
Tuesday
345914 01/18/11 15:59 fansler RQ3 N2
Monday
345913 01/17/11 17:57 David Anez Right HRS Electronics
345912 01/17/11 17:14 atrig Trigger-Download
345911 01/17/11 17:05 atrig Trigger-Download
345910 01/17/11 17:03 atrig Trigger-Download
345908 01/17/11 11:17 A. Camsonne Hall A BCM signal split and sent to HAPPEX electronics
Sunday
345907 01/16/11 14:27 atrig Trigger-Download
345906 01/16/11 14:27 atrig Trigger-Download
345905 01/16/11 14:24 atrig Trigger-Download
345904 01/16/11 14:23 atrig Trigger-Download
345903 01/16/11 14:20 atrig Trigger-Download
345902 01/16/11 14:19 atrig Trigger-Download
345901 01/16/11 14:17 atrig Trigger-Download
345900 01/16/11 14:16 atrig Trigger-Download
345899 01/16/11 14:14 atrig Trigger-Download
345898 01/16/11 14:13 atrig Trigger-Download
345897 01/16/11 14:10 atrig Trigger-Download
345896 01/16/11 14:06 atrig Trigger-Download
345895 01/16/11 11:40 fvansler LQ3 N2
Thursday
345894 01/13/11 17:28 A. Camsonne DVCS deadtime test runs
345893 01/13/11 17:27 A. Camsonne DVCS deadtime test runs
345890 01/13/11 14:03 atrig Trigger-Download
345888 01/13/11 12:23 A. Camsonne DVCS crates VME scaler
345887 01/13/11 11:52 A. Camsonne DVCS rac NIM crate
345886 01/13/11 11:47 A. Camsonne DVCS rack NIM crate
345885 01/13/11 11:46 A. Camsonne DVCS rack NIM crate
345884 01/13/11 09:20 fansler Cool down complete
345883 01/13/11 09:18 fansler RQ1 on cold return
Wednesday
345881 01/12/11 16:33 A. Camsonne Oscilloscope traces for run 9974 - adding 20 ms artificial dead time triggered by stop - S2m rate 45 Hz
345878 01/12/11 15:44 fansler Cool down
345877 01/12/11 15:41 A. Camsonne Oscilloscope traces for run 9973 - adding 10 ms artificial dead time triggered by stop - S2m rate 45 Hz
345874 01/12/11 14:54 A. Camsonne Oscilloscope traces for run 9972 - adding 1 ms artificial dead time triggered by stop - S2m rate 45 Hz
345871 01/12/11 14:20 fansler RD1 cold return
345870 01/12/11 13:56 A. Camsonne Oscilloscope traces for run 9970 - adding 100 us artificial dead time triggered by stop - S2m rate 45 Hz
345867 01/12/11 12:45 A. Camsonne Oscilloscope traces for run 9970 - adding 10 us artificial dead time triggered by stop - S2m rate 45 Hz
345866 01/12/11 12:43 A. Camsonne Oscilloscope traces for run 9970 - adding 10 us artificial dead time triggered by stop - S2m rate 45 Hz
345864 01/12/11 12:03 Fansler RQ3 on cold return
345857 01/12/11 09:47 A. Camsonne Oscilloscope traces for run 9966 - adding 1 us artificial dead time triggered by stop - S2m rate 450 Hz
345853 01/12/11 09:34 A. Camsonne Oscilloscope traces for run 9965 - adding 10 us artificial dead time triggered by stop - S2m rate 450 Hz
345850 01/12/11 09:25 A. Camsonne Oscilloscope traces for run 9964 - adding 100 us artificial dead time triggered by stop - S2m rate 450 Hz
345849 01/12/11 09:24 A. Camsonne Oscilloscope traces for run 9963 - adding 1 ms artificial dead time triggered by stop - S2m rate 450 Hz
345846 01/12/11 09:16 A. Camsonne Oscilloscope traces for run 9961 - adding 1 ms artificial dead time triggered by stop - S2m rate 450 Hz
345845 01/12/11 09:15 A. Camsonne Oscilloscope traces for run 9961 - adding 1 ms artificial dead time triggered by stop - S2m rate 50 Hz
345842 01/12/11 08:31 A. Camsonne Oscilloscope traces for run 9961 - adding 1 ms artificial dead time triggered by stop - S2m rate 50 Hz
345839 01/12/11 07:29 A. Camsonne Scalers run 9961
345838 01/12/11 07:16 A. Camsonne Oscilloscope traces for run 9961 - adding 20 ms artificial dead time triggered by stop - S2m rate 50 Hz
Tuesday
345835 01/11/11 19:04 Jack (Techs) VDC gas supply
345834 01/11/11 18:39 Jack (Techs) iocha16 (Right Arm magnets) is down
345833 01/11/11 18:31 A. Camsonne Oscilloscope traces for run 9960 - adding 20 ms artificial dead time pulsed rate 12.5 Hz - S2m rate 50 Hz
345830 01/11/11 16:23 A. Camsonne ROC14 2eSST VME traces
345829 01/11/11 16:15 A. Camsonne Oscilloscope traces for run 9959 - adding 20 ms artificial dead time pulsed rate 25 Hz - S2m rate 50 Hz
345827 01/11/11 16:00 A. Camsonne Oscilloscope traces for run 9958 - adding 20 ms artificial dead time pulsed rate 5 Hz - S2m rate 140 Hz
345823 01/11/11 15:35 A. Camsonne Oscilloscope traces for run 9957 - adding 20 ms artificial dead time pulsed rate 12.5 Hz - S2m rate 140 Hz
345822 01/11/11 15:34 A. Camsonne Oscilloscope traces for run 9957 - adding 20 ms artificial dead time pulsed rate 12.5 Hz - S2m rate 140 Hz
345819 01/11/11 12:00 A. Camsonne Oscilloscope traces for run 9956 - adding 20 ms artificial dead time pulsed rate 25 Hz - S2m rate 140 Hz
345816 01/11/11 11:50 A. Camsonne Oscilloscope traces for run 9955 - adding 20 ms artificial dead time pulsed rate 25 Hz - S2m rate 400 Hz
345815 01/11/11 11:48 A. Camsonne Oscilloscope traces for run 9955 - adding 20 ms artificial dead time pulsed rate 25 Hz - S2m rate 400 Hz
345812 01/11/11 11:33 A. Camsonne Trigger LED input unplugged
345810 01/11/11 11:20 A. Camsonne TS busy Run 9953
345809 01/11/11 11:19 A. Camsonne TS busy Run 9953
345807 01/11/11 10:37 A. Camsonne Oscilloscope traces for run 9953 - adding 20 ms uS artificial dead time self triggered at 40 Hz - S2m rate 440 Hz
345806 01/11/11 10:37 A. Camsonne Oscilloscope traces for run 9953 - adding 20 ms uS artificial dead time self triggered at 40 Hz - S2m rate 440 Hz
345805 01/11/11 10:36 A. Camsonne VME trace ROC17 in 2eSST
345800 01/11/11 10:20 A. Camsonne VME trace ROC17 in 2eSST
345794 01/11/11 09:50 A. Camsonne Bus analyzer moved to ROC 17 crate and old firmware loaded
345793 01/11/11 09:43 fansler Right dipole
345791 01/11/11 07:22 A. Camsonne Oscilloscope traces for run 9951 - adding 20 ms uS artificial dead time self triggered at 25 Hz - S2m rate 45 Hz
345788 01/11/11 06:47 A. Camsonne Oscilloscope traces for run 9949 - adding 20 ms uS artificial dead time self triggered at 25 Hz - S2m rate 400 Hz
345785 01/11/11 06:38 A. Camsonne Oscilloscope traces for run 9949 - adding 20 ms uS artificial dead time self triggered at 250 Hz - S2m rate 400 Hz
Monday
345777 01/10/11 18:19 A. Camsonne Oscilloscope traces for run 9944 - adding 10 ms artificial dead time correlated with trigger - S2m rate 400 Hz
345775 01/10/11 18:11 A. Camsonne Oscilloscope traces for run 9944 - adding 1.2 ms artificial dead time correlated with trigger - S2m rate 400 Hz
345772 01/10/11 17:48 A. Camsonne Oscilloscope traces for run 9943 - adding 800 uS artificial dead time correlated with trigger - S2m rate 400 Hz
345770 01/10/11 17:42 A. Camsonne Oscilloscope traces for run 9942 - adding 200 uS artificial dead time correlated with trigger - S2m rate 400 Hz
345769 01/10/11 17:42 A. Camsonne Oscilloscope traces for run 9942 - adding 200 uS artificial dead time correlated with trigger - S2m rate 400 Hz
345759 01/10/11 17:03 A. Camsonne Some S2m paddles turned on
345757 01/10/11 15:58 A. Camsonne Oscilloscope traces for run 9938 - adding 20 ms artificial dead time correlated with trigger
345754 01/10/11 15:09 A. Camsonne Oscilloscope traces for run 9937 - adding 50 ms artificial dead time correlated with trigger
345753 01/10/11 15:08 A. Camsonne Oscilloscope traces for run 9936 - adding 10 ms artificial dead time correlated with trigger
345752 01/10/11 15:08 A. Camsonne Oscilloscope traces for run 9936 - adding 10 ms artificial dead time correlated with trigger
345749 01/10/11 14:50 A. Camsonne hacweb1 was on but not on the network
345747 01/10/11 14:06 brads hacweb1 powered off?
345745 01/10/11 13:23 A. Camsonne Oscilloscope traces for run 9935 - adding 10 ms artificial dead time correlated with trigger
345742 01/10/11 12:01 A. Camsonne Oscilloscope traces for run 9935 - adding 8 ms artificial dead time correlated with trigger
345740 01/10/11 11:52 Fansler Right Arm cooldown
345738 01/10/11 10:54 A. Camsonne Oscilloscope traces for run 9934 - adding 2 ms artificial dead time correlated with trigger
345736 01/10/11 10:40 fansler Cooldown
345734 01/10/11 10:13 A. Camsonne Oscilloscope traces for run 9933 - adding 800 uS artificial dead time correlated with trigger
345732 01/10/11 09:58 A. Camsonne Oscilloscope traces for run 9932
345731 01/10/11 09:57 A. Camsonne Oscilloscope traces for run 9932
345730 01/10/11 09:54 A. Camsonne Oscilloscope traces for run 9931
345727 01/10/11 08:43 A. Camsonne Dead time test scaler rates run 9931
345726 01/10/11 08:38 A. Camsonne DVCS dead time test
Sunday
345721 01/09/11 11:37 V. Sulkosky re: Network issue with ROCs
Saturday
345720 01/08/11 23:17 brads re: Network issue with ROCs
345711 01/08/11 18:18 A. Camsonne Network issue with ROCs
Friday
Thursday
Wednesday
345676 01/05/11 10:41 R. Michaels ROC1 problem / solution
345674 01/05/11 10:04 fansler right arm cooldown
345673 01/05/11 10:02 A. Camsonne HV ON on Left HRS scintillators for DVCS dead time tests
345672 01/05/11 09:41 fansler Moller valves
Tuesday
345671 01/04/11 15:56 fansler Right arm valves
Monday
345670 01/03/11 14:36 A. Camsonne DVCS calorimeter unstacked
345669 01/03/11 12:34 Jack VDC high voltage is OFF
Sunday
345662 01/02/11 16:41 A. Camsonne DVCS IOC error message
Saturday
345660 01/01/11 10:35 A. Camsonne Calorimeter HV increased by 25 V